We have detected intermitted problems with the AS8 service were some accounts were not able to connect.
09:05 – Alerting reported problems with AS8. Engineers start to investigate.
09:25 – AS8 is restarted and moved to new hardware.
09:30 – Service restored.
09:35 – Service is still problematic. Engineers are working to resolve connectivity
09:45 – Problematic hardware was removed from pool and services have stabilised.
We apologies for any inconvenience.